# Self-Labeling the Job Shop Scheduling Problem

### v2

@@ 4.2 Self-Labeling Training Strategy
We propose a simple self-supervised training strategy that resorts only to the model being trained and does not require optimality information nor the formulation of the Markov Decision Process. Our strategy exploits two aspects: the ability of generative models to construct multiple (parallel) solutions, and the possibility of discriminating solutions based on their objective values, such as the makespan in the JSP. With these two ingredients, we design a procedure that at each iteration generates multiple solutions and uses the best one as a pseudo-label [30].
@@
More in detail, for each training instance $G$, we generate with the PN a set of $\beta$ different solutions. We do that by keeping $\beta$ partial solutions in parallel and independently sample for each one the next job from the probabilities generated by the decoder. Once the solutions have been generated, we take the one with the minimum makespan $\overline{\pi}$ and use it to compute the Self-Labeling loss ($\mathcal{L}_{\operatorname{SL}}$) by minimizing the cross-entropy on all the decision steps as follows:
$\mathcal{L}_{\operatorname{SL}}(\overline{\pi})=-\frac{1}{o}\sum_{t=1}^{o}\log f_{\theta}(y_{t}\,|\,\pi_{t},G),$ (5)
@@
where $o$ is the number of operations in the instance and $y_{t}\in J$ is the index of the job selected at the decision step $t$ while constructing the solution $\overline{\pi}$.
@@
The rationale of this training schema is to collect knowledge about sequences of decisions leading to high-quality solutions and save it in the model parameters $\theta$. This is done by selecting the best-generated solution $\overline{\pi}$ for an instance $G$ and maximizing its likelihood with Eq. 5 as in supervised learning, i.e., by minimizing cross-entropy. Through iterative exposure to different training instances, the model refines its ability to effectively solve the JSP.
@@
Although our strategy was developed independently, attentive readers may find a resemblance with the Cross-Entropy Method (CEM), a stochastic and derivative-free optimization method [15]. Typically applied to optimize parametric models such as Bernoulli and Gaussian mixtures models [52], the CEM relies on a maximum likelihood approach to either estimate a random variable or optimize the objective function of a problem [7]. According to Algorithm 2.2 in [7], the CEM independently samples $N=\beta$ solutions, selects a subset based on the problem’s objective ($\hat{\gamma}=S_{(N)}$ in our case), and updates the model’s parameters.
@@
Unlike the CEM, which independently tackles different instances by optimizing a separate model for each, our strategy trains a single model on many instances for learning how to globally solve the JSP. Moreover, we adopt a more complex parametric model rather than mixture models, always select a single solution for updating, and resorting to the gradient descent for updating parameters. Notably, our strategy also differs from CEM applications to model-based RL, e.g., [52], as we do not rely on rewards in Eq. 5. Summarizing, our strategy shares the idea of sampling and selecting with the CEM, but operates globally as a fully supervised learning once target solutions are identified. Hence, we named it: self-labeling training strategy.
@@
Other self-labeling approaches can be found e.g. in [10, 2], where the former uses K-Means to generate pseudo-labels, and the latter assigns labels to equally partition data with an optimal transportation problem. As highlighted in [2], these approaches may incur in degenerative solutions, i.e., solutions that trivially minimize Eq. 5, such as producing the same solution despite the input instance in our case. We remark that our strategy avoids such solutions by jointly using a probabilistic generation process and the objective value (makespan) of solutions to select the best pseudo-label $\overline{\pi}$.
@@ 5 Results
This section describes implementation details as well as selected competitors, and then presents the results. All the experiments were performed on an Ubuntu 22.04 machine equipped with an Intel Core i9-11900K and an NVIDIA GeForce RTX 3090 having 24GB of memory.
@@ 5.1 Experimental Setup
Architecture.
In all our experiments, we configure the PN of Section 4.1 in the same way. Our encoder consists of two GAT layers [8], both with 3 attention heads and Leaky slope at 0.15. In $\text{GAT}_{\texttt{1}}$, we set the size of each head to $64$ and concatenate their outputs; while in $\text{GAT}_{\texttt{2}}$, we increase the head’s size to $128$ and average their output to produce $e_{i}\in\mathbb{R}^{143}$ ($h=15+128$). Inside the decoder’s memory network, the MHA layer follows [46] except it concatenates the output of 3 heads with 64 neurons each, while $W_{\texttt{1}}\in\mathbb{R}^{11\times 192}$ and $W_{\texttt{2}}\in\mathbb{R}^{192\times 128}$ use 192 and 128 neurons, respectively. Thus, in Eq. 3, states $s_{j}\in\mathbb{R}^{d}$ have size $d=128$. Regarding the classifier, the FNN features a dense layer with 128 neurons activated through the Leaky-ReLU (slope $=0.15$) and a final linear with 1 neuron.
Training.
We train the PN with the strategy outlined in Section 4.2 on our dataset for $20$ epochs with the Adam optimizer [19] and learning rate at $0.0002$. During training, we save the parameters producing the lower average makespan on a hold-out set comprising 100 random instances of each shape. Additionally, we fix the number of sampled solutions $\beta$ in training to $256$, and use a batch size of 16, hence accumulating the gradient for 16 instances before updating $\theta$. Training this way roughly takes 120 hours, around 6 hours per epoch.
Competitors.
We compare the effectiveness of our methodology against various types of conventional and ML competitors reviewed in Section 2. We divide competitors as follows:
@@
  • Multiple (Randomized) constructives generate many solutions for an instance by introducing a controlled randomization in the selection process, a simple strategy for enhancing constructive algorithms [47, 25]. We consider the randomized results of CL and those of L2D. Since only greedy results are disclosed for L2D, we used the open-source code and trained model to generate randomized solutions as described at the bottom of Section 4.1. We also consider the three dispatching rules above, randomized by arbitrarily selecting the operation to schedule among the three with higher priority, and the Shifting Bottleneck Heuristic (SBH) [1] that creates multiple solutions while optimizing. All these approaches were seeded with 12345, generate $\beta=128$ solutions, and return the one with minimum makespan.
@@
For lack of space, we compare our PN against other learning proposals in Appendix B.
@@ 5.2 Performance on Benchmarks
Table 2: The average PGs of the algorithms on the two benchmark sets. For each row, we highlight in blue and bold the lowest (best) constructive and non-constructive gap, respectively. Shapes marked with * are larger than those seen in training by our SPN.
@@
SPT MWR MOR INSA L2D CL SPN
26.09 19.09 20.40 14.43 25.96 14.26 13.77
32.30 23.33 24.86 18.88 30.03 16.52 14.96
28.27 21.81 22.89 17.29 31.61 17.27 15.17
35.04 24.07 22.89 21.14 33.00 18.52 17.09
33.45 24.82 26.75 22.64 33.62 21.47 18.49
24.01 16.45 17.59 15.89 22.38 12.23 10.06
25.59 17.84 16.79 20.34 26.51 13.24 11.55
14.05 8.33 8.71 13.49 13.61 5.86 5.85
27.35 19.47 20.10 18.01 27.09 14.92 13.37
27.94 27.54 30.68 24.18 39.03 - 17.99
33.19 26.80 26.26 21.27 37.73 - 19.36
31.15 31.89 36.86 26.51 42.03 - 21.75
34.37 32.06 32.34 28.50 39.69 - 25.73
25.26 27.00 35.85 24.68 35.55 - 17.51
33.81 32.28 35.89 29.44 39.64 - 22.19
24.30 27.65 34.88 23.05 36.46 - 15.67
30.01 30.34 36.76 30.22 39.52 - 22.44
30.00 29.45 33.69 25.98 38.71 - 20.33
PDRs ($\beta=128$) RL ($\beta=128$) Our
SPT MWR MOR SBH L2D CL SPN128 SPN512
13.47 13.47 12.53 11.56 17.11 9.02 7.24 6.52
18.45 17.17 16.38 10.94 23.74 10.58 9.31 8.81
16.74 15.59 14.69 13.64 22.61 10.87 9.95 9.04
23.23 18.96 17.30 15.09 24.35 13.98 10.97 10.61
23.59 19.94 20.44 17.74 28.40 16.09 13.43 12.66
14.14 13.51 13.52 13.16 17.06 9.32 5.47 4.93
17.65 14.57 13.96 20.79 20.44 9.89 8.35 7.58
10.39 7.02 7.11 15.57 13.30 3.96 2.32 2.06
17.21 15.03 14.49 14.81 20.50 10.46 8.38 7.78
17.17 22.28 23.85 12.51 29.27 19.43 12.00 11.32
18.83 18.89 21.61 13.52 27.10 15.97 13.51 12.33
20.73 26.79 30.73 18.15 33.99 16.48 14.41 14.01
23.31 26.12 28.34 17.31 33.59 20.18 17.10 15.80
17.90 23.15 30.25 14.43 31.54 17.62 11.71 10.91
24.46 27.58 31.85 20.28 35.80 25.64 15.99 14.81
17.70 24.07 30.97 18.23 32.73 21.74 11.22 10.64
23.50 26.75 32.71 22.75 36.12 15.17 15.81 15.00
20.44 24.46 28.79 17.15 32.52 19.03 13.97 13.10
@@
This section evaluates the performance of our Self-labeling Pointer Network (SPN), configured and trained as explained in Section 5.1. When contrasted with greedy constructive approaches, our SPN generates a single solution by picking the job with the highest probability. In the other comparisons, we randomly sample 128 (SPN128) and 512 (SPN512) solutions as explained in Section 4.1.
We coded in Python 3.9 and PyTorch 13.1 our SPN, dispatching rules, INSA, SBH, MIP, and CP. While we report results from original papers of all the ML proposals but L2D, for which we used the open-source code for generating randomized solutions. Lastly, we evaluate performance on each benchmark instance in terms of the Percentage Gap: $\operatorname{PG}=100\cdot(C_{alg}\,/\,C_{ub}-1)$, where $C_{alg}$ is the makespan produced by an algorithm and $C_{ub}$ is either the instance optimal or best-known makespan.
@@
Table 2 presents the comparison of algorithms on Taillard’s and Demirkol’s benchmarks, each arranged in a distinct horizontal section. The table is vertically divided into Greedy Constructive, Multiple (Randomized) Constructive, and Non-constructive approaches; with the results of algorithms categorized accordingly. Each row reports the average PG on a specific instance shape while the last row (Avg) reports the average gap across all instances, regardless of their shapes.
@@
Table 2 shows that our SPN and CL produce lower gaps than PDRs, INSA, and SBH in both the greedy and randomized case, demonstrating the superiority of neural constructive approaches. Surprisingly, INSA and PDRs outperform L2D. This is likely related to how PDRs were coded in [49], where ours align with those in [25]. Focusing on our SPN, we observe that it consistently archives lower gaps than all the greedy approaches and, when applied in a randomized manner (SPN128), it outperforms all the randomized constructives. Notably, the SPN generalizes well to instance shapes larger than training ones (marked with * in Table 2), as its gaps do not progressively increase on such shapes.
@@
Our SPN is also competitive when compared with non-constructive approaches. The SPN512 largely outperforms the NLSA and L2S500 metaheuristics, and roughly align with L2S5k visiting 5000 solutions. On shapes larger than $20\times 20$, the SPN512 remarkably achieves lower gaps with just a few seconds of computations than a MIP executing for 3600 sec. Therefore, we conclude that our proposed SPN and training strategy are effective in solving the JSP.
@@
As already observed in [25, 50], our SPN and other neural constructive approaches are generally outperformed by CP and state-of-the-art metaheuristics, e.g., [35]. Although the performance gap is steadily narrowing, the higher complexity of CP solvers renders them more powerful, albeit at the cost of longer execution time (often dozens of minutes). Therefore, neural constructive approaches may be preferable whenever a quality solution must be provided in a few minutes.
@@
We refer the reader to Appendix D and Appendix E for statistical considerations and a comparison between our SPN and CP on very large instances, respectively.
6 Ablations
@@
This section studies additional aspects of our proposed methodology. For these evaluations, we adopt the same setting described in Section 5.1 and we explicitly indicate the varied parameters.
@@
Figure 2: SPN validation curves when trained with PPO and Self-Labeling in the same training setting of [49].

### v3

@@ 4.2 SLIM: Self-Labeling Improvement Method
We propose a self-supervised training strategy that uses the model’s output as a teaching signal, eliminating the need for optimality information or the formulation of Markov Decision Processes. Our strategy exploits two aspects: the capacity of generative models to construct multiple (parallel) solutions and the ability to discriminate solutions of combinatorial problems based on their objective values, such as the makespan in the JSP. With these two ingredients, we design a procedure that at each iteration generates multiple solutions and uses the best one as a pseudo-label [31].
@@
More in detail, for each training instance $G$, we randomly sample a set of $\beta$ solutions from the generative model $f_{\theta}(\cdot)$. We do that by keeping $\beta$ partial solutions in parallel and independently sample for each one the next job from the probabilities generated by the model. Once the solutions have been completely generated, we take the one with the minimum makespan $\overline{\pi}$ and use it to compute the Self-Labeling loss ($\mathcal{L}_{\operatorname{SL}}$). This loss minimizes the cross-entropy across all the steps as follows:
$\mathcal{L}_{\operatorname{SL}}(\overline{\pi})=-\frac{1}{|O|}\sum_{t=1}^{|O|}\log f_{\theta}(y_{t}\,|\,\overline{\pi}_{<t},G),$ (5)
@@
where $y_{t}\in J$ is the index of the job selected at the decision step $t$ while constructing the solution $\overline{\pi}$.
@@
The rationale behind this training schema is to collect knowledge about sequences of decisions leading to high-quality solutions and distill it in the parameters $\theta$ of the model. This is achieved by treating the best-generated solution $\overline{\pi}$ for an instance $G$ as a pseudo-label and maximizing its likelihood using Eq. 5. Similar to supervised learning, repeated exposure to various training instances enables the model to progressively refine its ability to solve combinatorial problems such as the JSP. Hence, we named this training strategy Self-Labeling Improvement Method (SLIM).
Relations with Existing Works
@@
Although our strategy was developed independently, attentive readers may find a resemblance with the Cross-Entropy Method (CEM), a stochastic and derivative-free optimization method [15]. Typically applied to optimize parametric models such as Bernoulli and Gaussian mixtures models [57], the CEM relies on a maximum likelihood approach to either estimate a random variable or optimize the objective function of a problem [7]. According to Algorithm 2.2 in [7], the CEM independently samples $N=\beta$ solutions, selects a subset based on the problem’s objective ($\hat{\gamma}=S_{(N)}$ in our case), and updates the parameters of the model.
@@
While the CEM independently tackles instances by optimizing a separate model for each, SLIM trains a single model on multiple instances to globally learn how to solve a combinatorial problem. Moreover, we adopt a more complex parametric model instead of mixture models, always select a single solution to update the model, and resort to the gradient descent for updating parameters $\theta$. Notably, our strategy also differs from CEM applications to model-based RL, e.g., [57], as we do not rely on rewards in Eq. 5. In summary, our strategy shares the idea of sampling and selecting with the CEM but globally operates as a supervised learning paradigm once the target solution is identified.
@@
Other self-labeling approaches can be found, e.g., in [10, 2], where the former uses K-Means to generate pseudo-labels, and the latter assigns labels to equally partition data with an optimal transportation problem. As highlighted in [2], these approaches may incur in degenerate solutions that trivially minimize Eq. 5, such as producing the same solution despite the input instance in our case. We remark that SLIM avoids such degenerate solutions by jointly using a probabilistic generation process and the objective value (makespan) of solutions to select the best pseudo-label $\overline{\pi}$.
@@ 5 Results
This section outlines the implementation details, introduces the selected competitors, and presents the key results. All the experiments were performed on an Ubuntu 22.04 machine equipped with an Intel Core i9-11900K and an NVIDIA GeForce RTX 3090 GPU having $24$ GB of memory. [^1] Our code is available at: https://github.com/AndreaCorsini1/SelfLabelingJobShop
@@ 5.1 Experimental Setup
Metric. We assess performance on each benchmark instance using the Percentage Gap: $\operatorname{PG}=100\cdot(C_{alg},/,C_{ub}-1)$, where $C_{alg}$ is the makespan produced by an algorithm and $C_{ub}$ is either the optimal or best-known makespan for the instance. Lower PG values indicate better results, as they reflect solutions with an objective value closer to the optimal or best-known makespan.
Architecture. In all our experiments, we configure the model of Sec. 4.1 in the same way. Our encoder consists of two GAT layers [8], both with $3$ attention heads and leaky slope at $0.15$. In $\text{GAT}_{\texttt{1}}$, we set the size of each head to $64$ and concatenate their outputs; while in $\text{GAT}_{\texttt{2}}$, we increase the head’s size to $128$ and average their output to produce $e_{i}\in\mathbb{R}^{143}$ ($h=15+128$). Inside the decoder’s memory network, the MHA layer follows [49] but it concatenates the output of $3$ heads with $64$ neurons each, while $W_{\texttt{1}}\in\mathbb{R}^{11\times 192}$ and $W_{\texttt{2}}\in\mathbb{R}^{192\times 128}$ use $192$ and $128$ neurons, respectively. Thus, the job states $s_{j}\in\mathbb{R}^{d}$ have size $d=128$. Finally, the classifier FNN features a dense layer with $128$ neurons activated through the Leaky-ReLU (slope $=0.15$) and a final linear with $1$ neuron.
Training. We train this generative model with SLIM (see Sec. 4.2) on our dataset for $20$ epochs, utilizing the Adam optimizer [19] with a constant learning rate of $0.0002$. In each training step, we accumulate gradients over a batch of size $16$, meaning that we update the model parameters $\theta$ after processing $16$ instances. During training and validation, we fix the number of sampled solutions $\beta$ to $256$ and save the parameters producing the lower average makespan on a hold-out set comprising $100$ random instances per shape included in our dataset. Training in this way takes approximately $120$ hours, with each epoch lasting around $6$ hours.
Competitors. We compare the effectiveness of our methodology against various types of conventional algorithms and ML competitors reviewed in Sec. 2. We divide competitors as follows:
@@
  • Multiple (Randomized) Constructives generate multiple solutions for an instance by introducing a controlled randomization in the selection process, a simple technique for enhancing constructive algorithms [50, 25]. We consider randomized results of CL and L2D. As only greedy results were disclosed for L2D, we used the open-source code to sample randomized solutions as described in Sec. 4.1. We also consider the three dispatching rules above, randomized by arbitrarily scheduling an operation among the three with higher priority, and the Shifting Bottleneck Heuristic (SBH) [1] that creates multiple solutions while optimizing. All these approaches were seeded with $12345$, generate $\beta=128$ solutions, and return the one with minimum makespan.
@@
For lack of space, we compare our methodology against other learning proposals in App. B.
@@ 5.2 Performance on Benchmarks
Table 2: The average PG of the algorithms on the benchmarks. In each row, we highlight in blue (bold) the best constructive (non-constructive) gap. Shapes marked with * are larger than those seen in training by our GM.
@@
SPT MWR MOR INSA L2D CL GM
26.1 19.1 20.4 14.4 26.0 14.3 13.8
32.3 23.3 24.9 18.9 30.0 16.5 15.0
28.3 21.8 22.9 17.3 31.6 17.3 15.2
35.0 24.1 22.9 21.1 33.0 18.5 17.1
33.4 24.8 26.8 22.6 33.6 21.5 18.5
24.0 16.4 17.6 15.9 22.4 12.2 10.1
25.6 17.8 16.8 20.3 26.5 13.2 11.6
14.0 8.3 8.7 13.5 13.6 5.9 5.8
27.4 19.5 20.1 18.0 27.1 14.9 13.4
27.9 27.5 30.7 24.2 39.0 - 18.0
33.2 26.8 26.3 21.3 37.7 - 19.4
31.2 31.9 36.9 26.5 42.0 - 21.8
34.4 32.1 32.3 28.5 39.7 - 25.7
25.3 27.0 35.8 24.7 35.6 - 17.5
33.8 32.3 35.9 29.4 39.6 - 22.2
24.3 27.6 34.9 23.0 36.5 - 15.7
30.0 30.3 36.8 30.2 39.5 - 22.4
30.0 29.4 33.7 26.0 38.7 - 20.3
PDRsβ=128 RLβ=128 Our
SPT MWR MOR SBH L2D CL GM128 GM512
13.5 13.5 12.5 11.6 17.1 9.0 7.2 6.5
18.4 17.2 16.4 10.9 23.7 10.6 9.3 8.8
16.7 15.6 14.7 13.6 22.6 10.9 10.0 9.0
23.2 19.0 17.3 15.1 24.4 14.0 11.0 10.6
23.6 19.9 20.4 17.7 28.4 16.1 13.4 12.7
14.1 13.5 13.5 13.2 17.1 9.3 5.5 4.9
17.6 14.6 14.0 20.8 20.4 9.9 8.4 7.6
10.4 7.0 7.1 15.6 13.3 4.0 2.3 2.1
17.2 15.0 14.5 14.8 20.5 10.5 8.4 7.8
17.2 22.3 23.8 12.5 29.3 19.4 12.0 11.3
18.8 18.9 21.6 13.5 27.1 16.0 13.5 12.3
20.7 26.8 30.7 18.2 34.0 16.5 14.4 14.0
23.3 26.1 28.3 17.3 33.6 20.2 17.1 15.8
17.9 23.2 30.2 14.4 31.5 17.6 11.7 10.9
24.5 27.6 31.8 20.3 35.8 25.6 16.0 14.8
17.7 24.1 31.0 18.2 32.7 21.7 11.2 10.6
23.5 26.8 32.7 22.8 36.1 15.2 15.8 15.0
20.4 24.5 28.8 17.2 32.5 19.0 14.0 13.1
@@
This section evaluates the performance of our generative Graph Model (GM), configured and trained with SLIM as explained in Sec. 5.1. When contrasted with greedy approaches, our GM generates a single solution by picking the job with the highest probability. In the other comparisons, we randomly sample $128$ (GM128) and $512$ (GM512) solutions as explained in Sec. 4.1. Note that we coded our GM, PDRs, INSA, SBH, MIP, and CP; while we reported results from original papers of all the ML competitors but L2D, for which we used the open-source code to generate randomized solutions.
@@
Tab. 2 presents the comparison of algorithms on Taillard’s and Demirkol’s benchmarks, each arranged in a distinct horizontal section. The table is vertically divided into Greedy Constructive, Multiple (Randomized) Constructive, and Non-constructive Approaches; with the results of algorithms categorized accordingly. Each row reports the average PG (the lower the better) on a specific instance shape while the last row (Avg) reports the average gap across all instances, regardless of their shapes.
@@
This table shows that our GM and CL produce lower gaps than PDRs, INSA, and SBH in both the greedy and randomized cases, proving the superiority of neural constructive approaches. Surprisingly, INSA and PDRs outperform L2D. This is likely related to how PDRs were coded in [54], where ours align with those in [25]. Focusing on our GM, we observe that it consistently archives lower gaps than all the greedy approaches and, when applied in a randomized manner (GM128), it outperforms all the randomized constructives. Notably, the GM generalizes well to instance shapes larger than training ones (marked with * in Tab. 2), as its gaps do not progressively increase on such shapes.
@@
Our GM is also competitive when compared with non-constructive approaches. The GM512 largely outperforms the NLSA and L2S500 metaheuristics, and roughly align with L2S5k visiting $5000$ solutions. On shapes larger than $20\times 20$, the GM512 remarkably achieves lower gaps with just a few seconds of computations than a MIP executing for $3600$ sec. Therefore, we conclude that our methodology, encompassing the proposed GM and SLIM, is effective in solving the JSP.
@@
As already observed in [25, 55], our GM and other neural constructive approaches are generally outperformed by CP and state-of-the-art metaheuristics, e.g., [36]. Although the performance gap is steadily narrowing, the higher complexity of CP solvers renders them more powerful, albeit at the cost of longer execution time (often dozens of minutes). Therefore, neural constructive approaches may be preferable whenever a quality solution must be provided in a few minutes.
@@
We refer the reader to Apps. E and D for statistical considerations and a comparison between our GM and CP on extremely large instances, respectively.
6 A Closer Look at SLIM
@@
This section studies additional aspects of our proposed methodology. For these evaluations, we adopt the same setting described in Sec. 5.1 and explicitly indicate the varied aspects and parameters.
@@
Figure 2: GM validation curves when trained with PPO and our SLIM in the same training setting of [54].